OpenStudy (anonymous):

subtract. simplify, if possible. 4x-36 over x^2-81 minus 4x-36 over 81-x^2

OpenStudy (anonymous):

\[\frac{4x-36}{x^2-81}-\frac{4x-36}{81-x^2}=\frac{4x-36}{x^2-81}+\frac{4x-36}{x^2-81}\] \[=\frac{8x-72}{x^2+81}\]}

OpenStudy (anonymous):

you have the correct setup, but not what they are looking for

OpenStudy (anonymous):

another typo, it should be \[=\frac{8x-72}{x^2-81}\] now maybe we can factor and cancel

OpenStudy (anonymous):

\[\frac{8(x-9)}{(x+9)(x-9)}=\frac{8}{x+9}\]

OpenStudy (anonymous):

thank you :) *muah*

OpenStudy (anonymous):

yw